OpenStudy (emilylizabeth):

Pre calculus help! Find the standard equation, center and radius. x^2+y^2+6y+9=1/16

OpenStudy (fifciol):

focus on this part: \[y^2 + 6y + 9\] can you write that in the form : \[(y - a)^2\]

OpenStudy (emilylizabeth):

Im not really sure how to do that.

OpenStudy (emilylizabeth):

could it be (y-3)^2?

OpenStudy (fifciol):

\[(y-3)^2=y^2-6y+9\] try \[(y+3)^2\]

OpenStudy (emilylizabeth):

y+6?

OpenStudy (fifciol):

\[(x-a)^2=x^2-2ax+a^2\] this is general formula. \[(x+3)^2=(x-(-3))^2+x^2-2x(-3)+(-3)^2=x^2+6x+9\]

OpenStudy (emilylizabeth):

So, all i'd do is solve it?

sam (.sam.):

Have you found the standard equation?

OpenStudy (emilylizabeth):

No, I haven't.

OpenStudy (fifciol):

@.Sam. this is not linear equation

sam (.sam.):

\[(x – h)^2 + (y – k)^2 = r^2?\]

OpenStudy (fifciol):

yup, that's it

sam (.sam.):

Teach her how to get to it

OpenStudy (fifciol):

I'm trying. Now that we found \[(y+3)^2=y^2+6y+9\], Can you substitute this into your original equation?

OpenStudy (emilylizabeth):

Substitute how?

OpenStudy (fifciol):

since your original equation is: \[x^2+y^2+6y+9=\frac{1}{16}\] and we've figured: \[y^2+6y+9=(y+3)^2\] we can now write down: \[x^2+(y+3)^2=\frac{1}{16}\]

OpenStudy (fifciol):

the standard formula has the following form: \[(x-a)^2+(y-b)^2=r^2\] wher (a,b) is the center and r is the radius

OpenStudy (fifciol):

so in our case, what is the center of a circle?

OpenStudy (emilylizabeth):

3 and x?

OpenStudy (fifciol):

we can write down our equation in slightly different form: \[(x-0)^2+(y-(-3))^2=\frac{1}{16}\]

OpenStudy (fifciol):

can you see it now?

OpenStudy (emilylizabeth):

0 and 3?

OpenStudy (fifciol):

0 and -3

OpenStudy (fifciol):

if we had (y-3)^2 it would be 3

OpenStudy (fifciol):

but we've got (y+3)^2 = (y-(-3))^2

OpenStudy (emilylizabeth):

So, how do we find the center?

OpenStudy (emilylizabeth):

Radius! Sorry.

OpenStudy (fifciol):

we can see that \[r^2=\frac{1}{16}\]

OpenStudy (fifciol):

thus\[r=\sqrt{\frac{1}{16}}\]

OpenStudy (fifciol):

so \[r = \frac{ \sqrt{1} }{ \sqrt{16}}=\frac{ 1 }{ 4 }\]

OpenStudy (emilylizabeth):

I think i understand it.

OpenStudy (fifciol):

glad to hear it :)

OpenStudy (emilylizabeth):

Thank you so much!

OpenStudy (mathmate):

@Fifciol It would be nice for her to start with the standard equation, so that she appreciates every step that you're showing her to do: As @.sam. said, the objective is: \(\large (x – h)^2 + (y – k)^2 = r^2\) It is much easier for a tutoree to follow you, and be motivated if he/she knows what the objective to start with. It also encourages him/her to try working out the problem, and to solve similar problems later on, which is the purpose of your lesson.

OpenStudy (fifciol):

right, maybe I'm not a good teacher :(